# High-Fidelity And Complex Test Data Generation For Real-World SQL Code Generation Services Industry Application & Experience Paper

### v1

@@ 3.4. Post-processor
3.5. Data validator
We would evaluate the generated data using a proto parser (Contributors, [n.d.]). This has an in-built check for the datatype and any other constraint on the proto fields. In case of any failure, we would be retrying the data generation, at most 2 retires, along with an error message. The integrated testing will be done by the framework.
@@ 3.5. Data validator
Figure 6. Semantic validation prompt template for generated test data. We also pass in the extracted SQL constraints, including the join information
We use the LLM for semantic validation (LLM-as-a-judge (Zheng et al., 2023)) and we pass the SQL constraints extracted from the pre-processor module along with the generated test data, as part of the template in Figure 6.
4. Evaluation results
@@ 4.1. Evaluation setup
4.2. Test coverage with more complex SQL query testing
@@
We assess the test coverage achieved by the generated test data. Our generation method can generate test data for complex schema with deeply nested structures. This facilitates more comprehensive testing, enabling more complex SQL test cases that were disabled due to the lack of such complex test data. Figure 7 presents improvement (%) in test coverage – we report the improvement instead of the raw test coverage values. The improvement in NL2SQL testing is more significant at $69.59\%$ because many available test cases were disabled due to restricted production data access. In SQL Code Assistant there are a lot more disabled test cases (mostly disabled because it is challenging to generate complex nested test data instances manually), and we expect to roll out new generated test data for those in the near future, which should increase the coverage further.
@@
Figure 7. Test coverage improvements (%) enabled by complex data generation and with more complex SQL test cases. NL2SQL test coverage improvement is much significant because many complex test cases were disabled due to restricted production table access before.
@@
4.3. Generation quality evaluation with complex nested data structures
@@
Next, we evaluate the completeness and correctness of the generation for complex nested data structures. Specifically we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types.
@@
Figure 8. Structural integrity & adherence to complex schema analysis. The scores are normalized (0 to 1) and the higher the better. We noticed some nested schema/proto definitions are inaccessible due to staleness and removal (invalid proto). Our data generation can generate test data compliant to complex nested schema structures from real-world scenarios.
@@
Figure 8 presents our structural integrity & adherence to complex data schema evaluation results. We use a subset of key evaluation criteria derived from the ones used in Table 2 – we excluded the criteria for primary key generation, which is always true for our test data. The evaluation scores are normalized (0 to 1) and the higher the better. Overall, our generated test data adheres very well to complex schema with deeply nested column types encountered in production SQL code generation services. There are some cases where nested field (proto) definitions are stale or removed, in which case the LLM guesses the field names and types for the inaccessible nested proto structures. Note that the scores are much higher (and near perfect) without such invalid proto definition cases w/o invalid proto. Note that the correct nesting levels scores for both with and without invalid proto are the same because we skip over those invalid nested structures in the aggregation – since we cannot assess the correct nesting levels for unknown proto. We typically see 3-level of nesting in complex SQL testing scenarios, and as deep as 5 levels.
4.4. A comparative study VS. production sample
@@
Figure 9. The number of distinct values generated for relevant predicates columns in the test SQL query and the natural language question, ”How many Shorts videos about Gaming were uploaded by (identified as) female creators in the United Kingdom (GB) during Q1 2024?”
@@
Figure 9 presents the number of distinct values for the (a) country codes column and (b) semantic tags column. Because the question is asking for the (identified as) female case, we also track the counts by (identified as) male and (identified as) female. In case of country codes, the production samples Prod include a more diverse set of countries than the Mock ones. This is OK because the question is looking for the answer in the United Kingdom (GB) specifically – and it is likely that our generation is less diverse because of this. At the same time, we also notice an imbalanced sampling between Prod/M and Prod/F in the production samples, which is not desirable as the question seeks to answer for the (identified as) female case. In case of semantic tags, which can entail different genres and keywords for different gaming contents (FPS, MMO, RPG, tutorial, review, etc.), our generated test data Mock does include a more diverse set of values. In both cases, the data is well balanced between the two select genders. Note that the test data is generated with syntax and semantic validations as described in Section 3.5, but the values can still be unrealistic as it is not learned from the actual production data (not available). Our generation approach also takes the test query/function into account to generate data that is more semantically aligned to the query. We discuss another example in Section 6.2.
5. Related works
@@
Test data generation for complex SQL testing: Test data generation is crucial for SQL query validation, query rewriting, automated query grading, and NL2SQL tasks. Common strategies involve using manually curated test data, production data sampling, constraints based data generation for table schema and SQL query (Somwase et al., 2024; Whiting et al., 2008; Bhangdiya et al., 2015; Chandra et al., 2015; Arcuri, 2019). Production data sampling can provide the most realistic test data for any complex schema, but it is often prohibited or infeasible in industrial contexts where service providers lack access to user data due to stringent privacy and security protections/policies. One of the key challenges in generating realistic high-fidelity test data is handling more complex nested schema structures and their inter-dependencies. Many SQL equivalence testing and constraints based SQL test data generation do not support arbitrarily nested structures and nested fields (array, struct and/or proto types) (He et al., 2024; Somwase et al., 2024) – limiting the test use cases to less complex SQL queries.
@@
Model-based tabular data generation: Despite the emerging interest in LLMs for test data generation (Gretel.ai, 2024; Lu et al., 2023; Karmarkar et al., 2024; Liu et al., 2023a), existing approaches often fail to fully address the complex testing requirements of industrial applications. While LLMs can provide realistic textual data and synthesize tabular data that adheres to schema format (Borisov et al., 2022), the out-of-the-box LLMs fall short at capturing subtle nuances and relationships betwen data atrributes (Xu et al., 2024). We also found that the vanilla LLM generation becomes more challenging and incomplete for a large nested schema (Table 2). Other deep generative model-based techniques have been used for synthetic tabular data (Zhang et al., 2023; Lee et al., 2023; Liu et al., 2023b). While producing realistic data examples that closely approximate the underlying distributions, the traditional tabular data generation solutions focus on flat schema without nested fields (struct, array, proto). Another challenge is to source good training datasets for learning the tabular generation model, since the production data is often restricted in the real-world testing environments. In this work, we propose a complex test data generation method based on a pre-trained LLM, gemini, and without any post-training.
6. Discussion
6.1. Scaling and cost of the generation
@@
While much more efficient and scalable than manual test data generation, as illustrated in Table 2, our test generation process can easily be deployed and parallelized across different batch of test queries/schemas, and also across different nested structures/fields within the same schema. The test data generation efficiency is important, also the main pain-point shared by the practitioners (e.g., developers submitting test queries and manual test data), as the data generation time takes much longer than the actual test generation/execution time.
@@
Table 3. Average ($\pm$ 1-std) number of accumulated input & output tokens, generation time per test data generation for the SQL code generation services. One can estimate the actual monetary cost using a (external) pricing model for gemini-2.0-flash.
Input tokens Output tokens Generation time (sec)
SQL Code Assistant 162199.5 ($\pm$ 350852.1) 3890 ($\pm$ 3121.5) 66.0 ($\pm$ 53.6)
NL2SQL 57082.3 ($\pm$ 102381.8) 8054.2 ($\pm$ 13040.1) 57.5 ($\pm$ 64.36)
@@
Table 3 presents a cost analysis of the proposed test data generation measured using the ”gemini-2.0-flash” model. It quantifies the average and standard deviation of the accumulated number of input/output tokens, as well as the generation time per test query. The high variability in token counts and generation time suggests a wide range of query and schema complexities. We report the number of input and output tokens instead of the monetary cost directly, as it depends on the pricing model (Google Cloud, 2024) (subject to change over time). The total cost of generation is proportional to the number of input/output tokens. Overall, our method can generate more comprehensive and complex test data (rows) much more efficiently than the manual generation – compared to the human labor cost, this is much cheaper and scalable, too.
6.2. Generating for test SQL query semantics
@@
Test data’s semantic coherence to its test target SQL query is critical in evaluating SQL code generation services, like NL2SQL. One of the key challenges in testing NL2SQL is evaluating semantic correctness of the predicted query (Floratou et al., 2024; Chung et al., 2025). While the NL2SQL prediction quality evaluation is beyond the scope of this study, a meaningful test data generation should provide a set of test data instances that are semantically relevant to the target SQL query. One way to address this is to cover all the edge cases for the relevant columns used in the SQL query, as we have focused in Section 4.4.
@@
Figure 10. Meaningful test data generation for this business analytics query requires generating data instances for each quarter in year 2022.
@@
6.3. Generation error analysis
@@
Figure 11. LLM hallucinates to introduce new columns. Only “fake_service_id”, “fake_service_dump_freshness_usec”, and (user) “id” are requested, but new fields are being introduced in the data.
@@
The last item is very difficult to fully address. We use LLM-as-a-judge to validate the semantic coherence of generated data, which runs the risk of false positives, but we do not have an oracle (or access to the real production tables) to ensure all relevant values are covered and generated. If we can provide some domain knowledge, as discussed in Section 6.2, that could help inform the LLM to generate for particular cases. This is an important problem to further improve the fidelity of mock data generation in the production settings. There are also techniques to statistically impute missing values and estimate potential impact of missing values on query results (Biessmann et al., 2019; Chung et al., 2018), which would be interesting to consider in the test data generation settings.
7. conclusion and future work
In this paper, we addressed the critical challenge of generating high-fidelity and complex test data for Google SQL code generation services, particularly in environments where production data access is restricted. We highlighted the limitations of traditional data generation methods, focusing on their inability to handle large, deeply nested schemas and maintain semantic integrity to the target test cases. By leveraging the latest gemini-2.0-flash model, coupled with strategic pre- and post-processing, we demonstrated the capability to generate comprehensive test data instances compliant to large complex schemas that effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries. Our experiments demonstrate the practical effectiveness of our approach, delivering a reliable method for generating realistic test data essential in industrial environments where access to production datasets is restricted. By deploying our test data generator, we’ve enabled more thorough SQL code testing, resulting in significantly improved test coverage.
@@
Acknowledgements.
We would like to thank John Cieslewicz, Kashmira Phalak and the anonymous reviewers for their valuable feedback. We also thank Yuchen Zhang for his contribution in the early stages of this project.

### v2

@@ III-D Post-processor
III-E Data validator
We would evaluate the generated data using a proto parser [12]. This has an in-built check for the datatype and any other constraint on the proto fields. In case of any failure, we would be retrying the data generation, at most 2 retires, along with an error message. The integrated testing will be done by the framework.
TABLE III: Enforced data validation rules. The generator extracts most of the syntactic and semantic constraints automatically, including the join paths. Additional input signals used for generation are also used for validation, like column correlation and constraints.
Rule Validation Requirements Validation Approach
$r_{1}$ No missing columns Use a proto parser to check for missing columns or proto fields.
$r_{2}$ Valid data types Use a proto parser to check for invalid data types. A schema column or proto field is invalid if its data type is different from the one specified in the fetched schema/proto.
$r_{3}$ Grouped column correlation Use statistical tools, like Pearson Correlation (numerical vs. numerical) and Chi-Square Test (categorical vs. categorical), to assert non-zero column correlations.
$r_{4}$ Join constraints Matches any join constraints extracted in the pre-processing. The column values in the joined tables at the join columns would be matched/copied from one table data to another.
$r_{5}$ Semantic constraints Semantic constraints are extracted from the test SQL query and they are validated using LLM-as-a-judge.
@@ III-E Data validator
Figure 8: Semantic validation prompt template for generated test data. We also pass in the extracted SQL constraints, including the join information and semantic constraints (e.g., relevant column values given the test query and question)
Semantic constraints and validation is not easy to capture with conventional approaches, and we use the LLM for semantic validation (LLM-as-a-judge [33]) and we pass the SQL constraints extracted from the pre-processor module along with the generated test data, as part of the template in Figure 8. Table III summarizes various syntactic and semantic data validation rules enforced by the generator; the generator retries the generation up to three times along with the error message upon any validation failure, including the semantic ones. In section IV-C, we evaluate how often the LLM hallucinates and returns semantically invalid data that passes the validation checks.
IV Evaluation results
@@ IV-A Evaluation setup
IV-B Test coverage with more complex SQL query testing
@@
We assess the test coverage achieved by the generated test data. Our generation method can generate test data for complex schema with deeply nested structures. This facilitates more comprehensive testing, enabling more complex SQL test cases that were disabled due to the lack of such complex test data. Table IV presents improvement (%) in test coverage and test accuracy – we report the improvement instead of the raw test coverage values. The improvement in NL2SQL testing coverage is more significant at $69.59\%$ because many available test cases were disabled due to restricted production data access. Notice that the generated test data also improves NL2SQL test accuracy. To our surprise, there were many cases where manually curated production test data that failed to differentiate predicted and golden SQL queries. For instance, some aggregate queries with wrong predicted filter predicates resulted in the correct answers (false positives); many test queries with INNER JOIN were flagged as incorrect (false negatives) against the manually curated test data, whereas adding sufficient test data for all the edge cases for the relevant columns used in the SQL query (Section VI-B) resulted in correct matching results. In SQL Code Assistant there are a lot more disabled test cases (mostly disabled because it is challenging to generate complex nested test data instances manually), and we expect to roll out new generated test data for those in the near future, which should increase the coverage further.
@@
TABLE IV: Test coverage and test accuracy changes (%) enabled by complex data generation and with more complex SQL test cases. NL2SQL test coverage improvement is much significant because many complex test cases were disabled due to restricted production table access before; NL2SQL test accuracy also improved significantly, since the manually curated test data allowed many false positives and false negatives.
Metric SQL Code Assistant NL2SQL
Test coverage (%) +12.58 +69.59
Test accuracy (%) +0.0 +20.4
@@
IV-C Generation quality evaluation with complex nested data structures
@@
Next, we evaluate the completeness and correctness of the generation for complex nested data structures. Specifically, we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types.
@@
Figure 9: Structural integrity & adherence to complex schema analysis. The scores are normalized (0 to 1) and the higher the better. We noticed some nested schema/proto definitions are inaccessible due to staleness and removal (invalid proto). Our data generation can generate test data compliant to complex nested schema structures from real-world scenarios.
@@
Figure 9 presents our structural integrity & adherence to complex data schema evaluation results. We use a subset of key evaluation criteria derived from the ones used in Table II – we excluded the criteria for primary key generation, which is always true for our test data. The evaluation scores are normalized (0 to 1) and the higher the better. Overall, our generated test data adheres very well to complex schema with deeply nested column types encountered in production SQL code generation services. There are some cases where nested field (proto) definitions are stale or removed, in which case the LLM guesses the field names and types for the inaccessible nested proto structures. Note that the scores are much higher (and near perfect) without such invalid proto definition cases w/o invalid proto. Note that the correct nesting levels scores for both with and without invalid proto are the same because we skip over those invalid nested structures in the aggregation – since we cannot assess the correct nesting levels for unknown proto. We typically see 3-level of nesting in complex SQL testing scenarios, and as deep as 5 levels.
IV-D Semantic data constraints validation with LLM-as-a-judge
The generator extracts semantic data constraints, like which column values are relevant to the given SQL query in the pre-processing step. We use the LLM as LLM-as-a-judge to validate the generated test data against the extracted semantic constraints. If some generated data instances fail to satisfy the constraints according to the LLM, then we retry with an additional hint that the previous generation failed to satisfy the failed semantic constraints. In all test cases, we have successfully generated test data instances after 1 or 2 retries. Interestingly, we observed that the LLM can hallucinate and actually pass invalid data instances as valid. Table V illustrates how generating multiple data instances per test case can limit the hallucination error rates, defined as the proportion of test cases with no semantically valid test data instances after $n$ instances generated. We report the hallucination error rates for NL2SQL test cases only because the unit test cases for SQL Code Assistant are too simple and contains simple semantic constraints, like COUNTIF(experiment_group = “GROUP A”) AS group_a_count. While most semantic constraint violations are resolved with a single additional data generation attempt ($n=2$), a few constraints remain challenging for the LLM-as-a-judge to validate. Specifically, complex datetime filter constraints that require type casting, time zone conversion, and date extraction consistently failed even after five attempts ($n=5$). For example, consider the constraint DATE(TIMESTAMP_SECONDS(CAST(status.time_processed _sec AS INT64)), ’America/Los_Angeles’) $\geq$ “2023-07-01”. Given a generated value like ”1678886400” for time_processed_sec, the validator correctly performed the initial conversion, recognizing that the timestamp corresponds to the date “2023-03-15”. However, it then failed to perform the final step of evaluating whether this date satisfied the constraint (i.e., that “2023-03-15” is not greater than or equal to “2023-07-01”). A potential remedy involves invoking a reasoning model or chain-of-thought prompting [34], though this significantly increases generation time. We plan to explore more cost-efficient ways to address these hallucination errors in future work.
TABLE V: LLM hallucination error rates for semantic data validation with repeated data generation ($n$, the number of test data instances) per test case. By generating multiple instances $n>1$, we get at least one test instance that satisfies the semantic constraints for most of the test cases. Semantic validation for SQL Code Assistant test cases was perfect, since the unit test cases did not involve any complex semantic constraints.
Number of instances per test ($n$)
1 2 3 4 5
NL2SQL 6.08% 2.02% 2.02% 2.02% 2.02%
IV-E A comparative study VS. production sample
@@ IV-D Semantic data constraints validation with LLM-as-a-judge
Figure 10: The number of distinct values generated for relevant predicates columns in the test SQL query and the natural language question, ”How many Shorts videos about Gaming were uploaded by (identified as) female creators in the United Kingdom (GB) during Q1 2024?”
@@
Figure 10 presents the number of distinct values for the (a) country codes column and (b) semantic tags column. Because the question is asking for the (identified as) female case, we also track the counts by (identified as) male and (identified as) female. In case of country codes, the production samples Prod include a more diverse set of countries than the Mock ones. This is OK because the question is looking for the answer in the United Kingdom (GB) specifically – and it is likely that our generation is less diverse because of this. At the same time, we also notice an imbalanced sampling between Prod/M and Prod/F in the production samples, which is not desirable as the question seeks to answer for the (identified as) female case. In case of semantic tags, which can entail different genres and keywords for different gaming contents (FPS, MMO, RPG, tutorial, review, etc.), our generated test data Mock does include a more diverse set of values. In both cases, the data is well balanced between the two select genders. Note that the test data is generated with syntax and semantic validations as described in Section III-E, but the values can still be unrealistic as it is not learned from the actual production data (not available). Our generation approach also takes the test query/function into account to generate data that is more semantically aligned to the query. We discuss another example in Section VI-B.
V Related works
@@
Test data generation for complex SQL testing: Test data generation is crucial for SQL query validation, query rewriting, automated query grading, and NL2SQL tasks. Common strategies involve using manually curated test data, production data sampling, constraints based data generation for table schema and SQL query [5, 1, 3, 4, 6]. Production data sampling can provide the most realistic test data for any complex schema, but it is often prohibited or infeasible in industrial contexts where service providers lack access to user data due to stringent privacy and security protections/policies. One of the key challenges in generating realistic high-fidelity test data is handling more complex nested schema structures and their inter-dependencies. Many SQL equivalence testing and constraints based SQL test data generation do not support arbitrarily nested structures and nested fields (array, struct and/or proto types) [11, 5] – limiting the test use cases to less complex SQL queries.
@@
Model-based tabular data generation: Despite the emerging interest in LLMs for test data generation [16, 19, 20, 21], existing approaches often fail to fully address the complex testing requirements of industrial applications. While LLMs can provide realistic textual data and synthesize tabular data that adheres to schema format [27], the out-of-the-box LLMs fall short at capturing subtle nuances and relationships betwen data atrributes [25]. We also found that the vanilla LLM generation becomes more challenging and incomplete for a large nested schema (Table II). Other deep generative model-based techniques have been used for synthetic tabular data [29, 30, 31]. While producing realistic data examples that closely approximate the underlying distributions, the traditional tabular data generation solutions focus on flat schema without nested fields (struct, array, proto). Another challenge is to source good training datasets for learning the tabular generation model, since the production data is often restricted in the real-world testing environments. In this work, we propose a complex test data generation method based on a pre-trained LLM, gemini, and without any post-training.
VI Discussion
VI-A Scaling and cost of the generation
@@
While much more efficient and scalable than manual test data generation, as illustrated in Table II, our test generation process can easily be deployed and parallelized across different batch of test queries/schemas, and also across different nested structures/fields within the same schema. The test data generation efficiency is important, also the main pain-point shared by the practitioners (e.g., developers submitting test queries and manual test data), as the data generation time takes much longer than the actual test generation/execution time.
@@
TABLE VI: Average ($\pm$ 1-std) number of accumulated input & output tokens, generation time per test data generation for the SQL code generation services. One can estimate the actual monetary cost using a (external) pricing model for gemini-2.0-flash.
Metric SQL Code Assistant NL2SQL
Input tokens 162199.5 ($\pm$ 350852.1) 57082.3 ($\pm$ 102381.8)
Output tokens 3890 ($\pm$ 3121.5) 8054.2 ($\pm$ 13040.1)
Generation time (sec) 66.0 ($\pm$ 53.6) 57.5 ($\pm$ 64.36)
@@
Table VI presents a cost analysis of the proposed test data generation measured using the ”gemini-2.0-flash” model. It quantifies the average and standard deviation of the accumulated number of input/output tokens, as well as the generation time per test query. The high variability in token counts and generation time suggests a wide range of query and schema complexities. We report the number of input and output tokens instead of the monetary cost directly, as it depends on the pricing model [35] (subject to change over time). The total cost of generation is proportional to the number of input/output tokens. Overall, our method can generate more comprehensive and complex test data (rows) much more efficiently than the manual generation – compared to the human labor cost, this is much cheaper and scalable, too.
VI-B Generating for test SQL query semantics
@@
Test data’s semantic coherence to its test target SQL query is critical in evaluating SQL code generation services, like NL2SQL. One of the key challenges in testing NL2SQL is evaluating semantic correctness of the predicted query [36, 37]. While the NL2SQL prediction quality evaluation is beyond the scope of this study, a meaningful test data generation should provide a set of test data instances that are semantically relevant to the target SQL query. One way to address this is to cover all the edge cases for the relevant columns used in the SQL query, as we have focused in Section IV-E.
@@
Figure 11: Meaningful test data generation for this business analytics query requires generating data instances for each quarter in year 2022.
@@
VI-C Generation error analysis
@@
Figure 12: LLM hallucinates to introduce new columns. Only “fake_service_id”, “fake_service_dump_freshness_usec”, and (user) “id” are requested, but new fields are being introduced in the data.
@@
The last item is very difficult to fully address. We use LLM-as-a-judge to validate the semantic coherence of generated data, which runs the risk of false positives, but we do not have an oracle (or access to the real production tables) to ensure all relevant values are covered and generated. If we can provide some domain knowledge, as discussed in Section VI-B, that could help inform the LLM to generate for particular cases. This is an important problem to further improve the fidelity of mock data generation in the production settings. There are also techniques to statistically impute missing values and estimate potential impact of missing values on query results [38, 39], which would be interesting to consider in the test data generation settings.
VII conclusion and future work
In this paper, we addressed the critical challenge of generating high-fidelity and complex test data for Google SQL code generation services, particularly in environments where production data access is restricted. We highlighted the limitations of traditional data generation methods, focusing on their inability to handle large, deeply nested schemas and maintain semantic integrity to the target test cases. By leveraging the latest gemini-2.0-flash model, coupled with strategic pre- and post-processing, we demonstrated the capability to generate comprehensive test data instances compliant to large complex schemas that effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries. Our experiments demonstrate the practical effectiveness of our approach, delivering a reliable method for generating realistic test data essential in industrial environments where access to production datasets is restricted. While our implementation targeted Google SQL, the fundamental methodology is dialect-agnostic and should be readily applicable to other SQL code generation services. By deploying our test data generator, we’ve enabled more thorough SQL code testing, resulting in significantly improved test coverage.
@@
Acknowledgment
We would like to thank John Cieslewicz, Kashmira Phalak, Yannis Papakonstantinou and the anonymous reviewers for their valuable feedback. We also thank Yuchen Zhang for his contribution in the early stages of this project.